OpenStudy (anonymous):

lim as x approaches 3 of (1/x-3)(1/sqrt(x+1)-1/2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} \frac{ 1 }{ x-3 }(\frac{ 1 }{ \sqrt{x+1} }-\frac{ 1 }{ 2 })\]

OpenStudy (fibonaccichick666):

so what type of indeterminant to you get?

OpenStudy (fibonaccichick666):

to start, so we can see if it is something we can work with :)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I just know that you're supposed to multiply by the conjugate.

ganeshie8 (ganeshie8):

Fib is asking if something bad happens when you plugin 3 into the expression

OpenStudy (fibonaccichick666):

that is what I am asking essentially^ :)

OpenStudy (anonymous):

0/0

OpenStudy (fibonaccichick666):

how are you getting a zero on top? but besides the point

OpenStudy (anonymous):

I just put that, but it comes out to just zero.

OpenStudy (anonymous):

It's 0/0

OpenStudy (anonymous):

When you plug in the 3, you get 1/0 (0) so it's 0/0, isn't it?

OpenStudy (anonymous):

I just need help getting passed the first step.

OpenStudy (anonymous):

Yes, it is \(\frac 00\).

OpenStudy (fibonaccichick666):

no, 1/0 is just 1/0.

OpenStudy (anonymous):

But times the zero inside of the parenthesis?

OpenStudy (fibonaccichick666):

yes it is times zero, but that is an inderterminant form, times zero. so we can't just say that as eaily

OpenStudy (anonymous):

Isn't this a little beside the point, though?

OpenStudy (fibonaccichick666):

so personally I would use squeeze theorem here, have you learned it?

OpenStudy (anonymous):

\(1/0\) is not an indeterminate form.

OpenStudy (anonymous):

And \(1\times 0 = 0\)

OpenStudy (fibonaccichick666):

1/0 is inderterminant. 0/0 is underfined

OpenStudy (fibonaccichick666):

crap got em backwards yet again didn't I?

OpenStudy (anonymous):

...

OpenStudy (fibonaccichick666):

ugh. I always do that, but doesn't really matter, so anyways, have you learned squeeze theorem?

OpenStudy (anonymous):

Okay, so, now that that tangent has been gotten out of our systems :)

OpenStudy (anonymous):

Maybe I have, but I've never heard of it by that name? I'm not sure.

OpenStudy (fibonaccichick666):

it can also be called sandwich or something like that

OpenStudy (fibonaccichick666):

if not we can find another way I'm sure

OpenStudy (anonymous):

Can you maybe just show me the first step? Because I did the conjugate and got x-1. I came out with the correct answer, but the first step is marked wrong.

OpenStudy (fibonaccichick666):

ok, so what conjugate did you use?

OpenStudy (fibonaccichick666):

show me that step

OpenStudy (anonymous):

I used \[\sqrt{x-1}\]

OpenStudy (fibonaccichick666):

ok can you show me how you did it? It might just be a quick fix then

OpenStudy (anonymous):

I think we can do this with algebra

OpenStudy (anonymous):

One thing I like to do is re-parametrize... so I let \(u=x-3\) and \(x=u+3\), because \(0\) limits seem easier to work with.

OpenStudy (fibonaccichick666):

that is true, but wait a sec, he is showing his steps. Although I do think squeeze is the easiest once learned

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} \frac{ 1 }{ x-3 }(\frac{ 1 }{ \sqrt{x+1} }-\frac{ 1 }{ 2 })\] Becomes: \[ \lim_{u\to 0}\frac{1}{u}\left(\frac{ 1 }{ \sqrt{u+4} }-\frac{ 1 }{ 2 }\right) \]

OpenStudy (anonymous):

\[(\frac{ 1 }{ \sqrt{x+1} }-\frac{ 1 }{ 2 }) \times \frac{ \sqrt{x-1} }{ \sqrt{x-1} } = \frac{ \sqrt{x-1} }{ x-1 } - \frac{ \sqrt{x-1} }{ 2 \times \sqrt{x-1} }\]

OpenStudy (anonymous):

No, conjugates don't work that way.

OpenStudy (anonymous):

That's what my teacher said. My question is why.

OpenStudy (anonymous):

The conjugate of \(\sqrt{x+1}\) is just \(\sqrt{x+1}\). However, the conjugate of \(\sqrt{a}+b\) is \(\sqrt{a}-b\).

OpenStudy (anonymous):

Ah. Okay.

OpenStudy (fibonaccichick666):

yea, what he said, got there before I could

OpenStudy (triciaal):

@FibonacciChick666 what is the name for what you are referring to as the "squeeze theorem"?

OpenStudy (fibonaccichick666):

squeeze theorem is it's name. It is the most common name used

OpenStudy (anonymous):

You guys are just heading in all types of directions.

OpenStudy (anonymous):

Maybe reign it in a little bit.

OpenStudy (fibonaccichick666):

no, but wio is right about that step.

OpenStudy (anonymous):

Yes, I see that. I know this. However, what is the right way to go about it?

OpenStudy (fibonaccichick666):

although I was extremely happy to see you did multiply by a fraction equivalent to 1

OpenStudy (anonymous):

I think it will be less confusing if you start with this:\[ \lim_{u\to 0}\frac{1}{u}\left(\frac{ 1 }{ \sqrt{u+4} }-\frac{ 1 }{ 2 }\right) \]

OpenStudy (anonymous):

....

OpenStudy (fibonaccichick666):

^that too.

OpenStudy (anonymous):

And the thing I would want to do first is get rid of fractions by multiplying by the lcm. In this case it would be \(2\sqrt{u+4}\).

OpenStudy (fibonaccichick666):

I agree,

OpenStudy (anonymous):

Okay. Never mind. Someone else showed me the right way to do it.

OpenStudy (fibonaccichick666):

there are like 10 ways to go about this

OpenStudy (fibonaccichick666):

not just one correct way

OpenStudy (anonymous):

"the" right way, lol

OpenStudy (anonymous):

And I just needed 1.

OpenStudy (anonymous):

Wait, how many people are helping you with this problem at the same time?

OpenStudy (anonymous):

Too many, apparently lol. But thanks for the generosity in trying to help.

OpenStudy (fibonaccichick666):

well, I tried giving you one, you didn't care for it. Now wio is giving you another one, which is a very nice way, and you are disregarding it.

OpenStudy (anonymous):

I'm not disregarding it.

OpenStudy (anonymous):

Please, try not to misinterpret my intentions.

OpenStudy (triciaal):

@wio do you have any idea what he means by the "squeeze theorem?"

OpenStudy (fibonaccichick666):

Wio's method he is about to give is probably the most elegant algebraic way

OpenStudy (anonymous):

I don't need elegant. I just needed a direction to go in, and you guys took the most roundabout way of getting there.

OpenStudy (anonymous):

Squeeze theorem is \[ g(x)\leq f(x)\leq h(x) \] where \[ \lim g(x)=L,\quad \lim h(x) = L \]

OpenStudy (anonymous):

I'm not disregarding anything. I just wanted an explanation as to why my way was incorrect. You guys gave it to me. However, more questions were added into the mix than necessary.

OpenStudy (fibonaccichick666):

not intentionally lol, but after all, math happens :)

OpenStudy (anonymous):

The people on this site used to be a lot nicer and a lot more focused.

OpenStudy (fibonaccichick666):

and anyways, I think you should look at wio's method. It is much simpler than the way you were going about it.

OpenStudy (zarkon):

let \(f(x)=\dfrac{1}{\sqrt{x+1}}\) then \[\lim_{x \to3} \frac{ 1 }{ x-3 }(\frac{ 1 }{ \sqrt{x+1} }-\frac{ 1 }{ 2 })=\lim_{x\to3}\frac{f(x)-f(3)}{x-3}=f'(3)\]

OpenStudy (anonymous):

I suppose that does work when you have \(0/0\) forms, hmmm

OpenStudy (fibonaccichick666):

@Zarkon love that, i din't even see that relation too

OpenStudy (anonymous):

Thank you, Zarkon. This is the way that I'm used to. Thank you. This is how people should answer questions. Just quick and to the point.

OpenStudy (fibonaccichick666):

no, that is just giving an answer and not what this site is about. He only gave it to show yet another way of solving this problem

OpenStudy (anonymous):

@jayy2014 We don't even know that you know the derivative though...

OpenStudy (fibonaccichick666):

if you don't know squeeze theorem, you don't know derivative

OpenStudy (anonymous):

I can find it. I'm not new to Calculus. I just needed a refresher.

OpenStudy (anonymous):

This site isn't about giving answers; I know this. I'm not new here.

OpenStudy (fibonaccichick666):

which is why, you had 3 views and 3 ways to go about this problem

OpenStudy (anonymous):

But you guys were taking too long to tell me about those three ways.

OpenStudy (anonymous):

You were responding to everyone at the same time. It's like a teacher trying to answer three questions at once. Someone is going to leave unsatisfied.

OpenStudy (fibonaccichick666):

... this is math not McDonald's. It takes time to learn and understand the material. And with group answering that happens. then listen to wio, he has a more streamlined style

OpenStudy (anonymous):

Either way, I see where I made my mistake, and it was a pretty simple one. I'm surprised at myself. Thanks for the help, as unorganized as it was. I'm done with this problem, and this conversation.

OpenStudy (anonymous):

Math, not McDonald's. Pfft. I'm done.

OpenStudy (triciaal):

@wio and @Zarkon thank you @jayy2014 sorry if I invaded your space

hartnn (hartnn):

 Though Zarkon's Method is elegant, the first though which comes to mind while seeing something in sqrt is to rationalize :)